OpenStudy (anonymous):

What is arccos(-1/6)? Please explain how to solve it without a calculator. Thanks!

OpenStudy (anonymous):

I don't know. But what if you graphed it and figured it out from that?

OpenStudy (anonymous):

Yes, I understand how to solve it using graphing or a calculator. The full problem is: tan[arccos(-1/6)] which I had to solve by hand. Like, for: arcsin[sin(5pi/4)] I can easily just flip it around because it contains a pi. This one doesn't though, so I'm a bit confused.

OpenStudy (anonymous):

You know more about it than I do--right now I did find this: Since none of the six trigonometric functions are one-to-one, they are restricted in order to have inverse functions. Therefore the ranges of the inverse functions are proper subsets of the domains of the original functions So, that gives me a hint.

OpenStudy (anonymous):

Just asking: doesn't that mean you have to turn one-sixth into radians?

OpenStudy (wolf1728):

Cosine 45° = square root of 2 The arc cosine of (1/6) = 80.4059317731 degrees = 1.4033482476 radians

OpenStudy (anonymous):

Well, for arcsin[sin(5pi/4)]. sin(5pi/4) = \[-\sqrt{2}/2\] Then, arcsin\[-\sqrt{2}/2\] would be \[-\sqrt{\pi}/4\]

OpenStudy (anonymous):

I'm looking for the exact values. P:. Oops, forgot to mention that.

OpenStudy (wolf1728):

I'd say it doesn't have to bee converted to radians, since it is a ratio of two sides and not an angle.

OpenStudy (anonymous):

Good point

OpenStudy (anonymous):

BUT, if it has to be an "exact value" she does have to use the square root in this case, right?

OpenStudy (anonymous):

Yes, it can have a square root.

OpenStudy (anonymous):

I think it has to have a square root. So, wouldn't you just graph it with the square root location, just as you would a specific number. Just curious.

OpenStudy (wolf1728):

Well the arc tangent of sqrt(3) = 60° (2 exact values)

OpenStudy (anonymous):

Ah, I see. Thanks.

OpenStudy (anonymous):

This was just a really odd question, but I think I've got it now. Thank you guys. :).

OpenStudy (anonymous):

:)

OpenStudy (wolf1728):

okay :-)